A time varying net force acting on a4.0kg particle causes the particle to have a displacement given by x=2.0t-3.0t2+1.0t3wher x in meters t in seconds

1
Expert's answer
2021-12-12T16:43:27-0500

I assume the task is to find the force. Let "m=4kg" be the mass of the particle. According to the second Newton's law, the net force on the particle is:


"F =ma"

where "a" is the acceleration of the particle. By definition, the acceleration is the second derivative from the coordinate: "a(t) = x''(t)" . Since the coordinate is given as follows:


"x(t) = 2t-3t^2+t^3"

obtain the acceleration:


"a(t) = (2t-3t^2+t^3)'' = 6t-6"

Thus, the force is:


"F(t) = 4\cdot (6t-6)=24t-24\space\space (N)"

Answer. "F(t)=24t-24\space\space (N)"